OpenStudy (anonymous):

Find the polynomial function with roots 11 and 2i. help please:)

OpenStudy (anonymous):

(x-11)(x-2i)=x^2-(11+2i)x+22i?

OpenStudy (anonymous):

if 2i is a root then -2i must be root

OpenStudy (anonymous):

(x-11)(x-2i)(x+2i)=0

OpenStudy (anonymous):

or for a real polynomial (x-11)(x^2+4)=x^3-11x^2+4x-44 but this also has -2i as a root as cinar says

OpenStudy (anonymous):

Thank you:)